OpenStudy (anonymous):

Complete this statement: A polygon with all sides the same length is said to be ______. (1 point)regular equilateral equiangular convex

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

equilateral means all the sides are congruent, equiangular mean all the angles are congruent(:

OpenStudy (anonymous):

Complete this statement: A polygon with all sides the same length is said to be ______. (1 point)regular equilateral equiangular convex NOTE: Question 2 was entered in error. You will receive credit for ANY answer. 2. A road sign is in the shape of a regular heptagon. What is the measure of each angle on the sign? Round to the nearest tenth. (1 point)128.6o 900o 64.3o 231.4o 3. In the figure, the horizontal lines are parallel and AB = BC = CD. Find the measure of JM. The diagram is not drawn to scale. (1 point)14 21 28 7 4. Short Answer Note: Remember to show all of the steps that you use to solve the problem for questions 4–10. You can use the comments field to explain your work. Your teacher will review each step of your responses. Find the sum of the measures of the interior angles of the figure. (1 point) 5. Find the missing values of the variables. The diagram is not drawn to scale. (3 points) 6. The sum of the measures of two exterior angles of a triangle is 264o. What is the measure of the third exterior angle? (3 points) 7. How many sides does a regular polygon have if each exterior angle measures 72o? (2 points) 8. Find the values of the variables x,y, and z in the parallelogram. The diagram is not drawn to scale. (4 points) 9. In parallelogram DEFG, DH = x + 1, HF = 3y, GH = 3x – 4, and HE = 5y + 1. Find the values of x and y. The diagram is not drawn to scale. (4 points) 10. If ON = 8x – 8, LM = 7x + 4, NM = x – 5, and OL = 3y – 6, find the values of x and y for which LMNO must be a parallelogram. The diagram is not drawn to scale. (3 points) Essay Your teacher will grade your response for questions 11–13 to ensure that you receive proper credit for your answers. 11. What are the missing reasons in the proof? Given: with diagonal Prove: (4 points) 12. Based on the information in the diagram, can you prove that the figure is a parallelogram? Explain. (4 points) 13. Given: and Prove: VUTS is a parallelogram (4 points)